OpenStudy (anonymous):

((x)/(11)-1)/(x-11) I keep getting x/11 as answer but its not correct

OpenStudy (pfenn1):

\[\left( x \over 11 \right)-1 \over x-11\] Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Is the answer (1/11)?

OpenStudy (anonymous):

yes it is thank you ;)

OpenStudy (pfenn1):

I ask, because I was checking to see if I solved the equation properly. I took the numerator and converted both terms to fractions \[\left( x \over 11 \right) -1=\left( x \over 11 \right)-\left( 11\over 11 \right)=\left( x-11 \over 11 \right)\]

OpenStudy (anonymous):

okay

OpenStudy (pfenn1):

Can you solve simplify the problem from here?

OpenStudy (anonymous):

yes I guess I was just confused with why it wasnt x but i see why it is 1/11

OpenStudy (pfenn1):

good.